# GPT4RoI: Instruction Tuning Large Language Model on Region-of-Interest

### v4

@@ Abstract
GPT4RoI: Instruction Tuning Large Language Model on Region-of-Interest
Shilong Zhang1,2,3thanks: Equal contribution.  Peize Sun1∗  Shoufa Chen1∗  Min Xiao2
Wenqi Shao2  Wenwei Zhang2  Yu Liu3  Kai Chen2  Ping Luo1,2
1The University of Hong Kong   2Shanghai AI Laboratory   3Alibaba Group
@@ 1 Introduction
Figure 1: Comparison of visual instruction tuning on image-text pairs and spatial instruction tuning on region-text pairs. The bounding box and text description of each object are provided in region-text datasets. During training, the bounding box is from annotations, and in inference, it can be provided by user or any off-the-shelf object detector
@@
Recent advancements of large language models (LLM) have shown incredible performance in solving natural language processing tasks in a human-like conversational manner, for example, commercial products  OpenAI (2022); Anthropic (2023); Google (2023); OpenAI (2023) and community open-source projects Touvron et al. (2023a, b); Taori et al. (2023); Chiang et al. (2023); Du et al. (2022); Sun and Xipeng (2022). Their unprecedented capabilities present a promising path toward general-purpose artificial intelligence models. Witnessing the power of LLM, the field of multimodal models Yang et al. (2023c); Huang et al. (2023); Girdhar et al. (2023); Driess et al. (2023) is developing a new technology direction to leverage LLM as the universal interface to build general-purpose models, where the feature space of a specific task is tuned to be aligned with the feature space of pre-trained language models.
@@
As one of the representative tasks, vision-and-language models align the vision encoder feature to LLM by instruction tuning on image-text pairs, such as MiniGPT-4 Zhu et al. (2023), LLaVA Liu et al. (2023a), InstructBLIP Dai et al. (2023), etc. Although these works achieve amazing multimodal abilities, their alignments are only on image-text pairs Chen et al. (2015); Sharma et al. (2018); Changpinyo et al. (2021); Ordonez et al. (2011); Schuhmann et al. (2021), the lack of region-level alignment limits their advancements to more fine-grained understanding tasks such as region caption Krishna et al. (2017) and reasoning Zellers et al. (2019a). To enable region-level understanding in vision-language models, some works attempt to leverage external vision models, for example, MM-REACT Yang et al. (2023c), InternGPT Liu et al. (2023d) and DetGPT Pi et al. (2023). However, their non-end-to-end architecture is a sub-optimal choice for general-purpose multi-modal models.
@@
Considering the limitations of previous works, our objective is to construct an end-to-end vision-language model that supports a fine-grained understanding on region-of-interest. Since there is no operation that can refer to specific regions in current image-level vision-language models Zhu et al. (2023); Liu et al. (2023a); Zhang et al. (2023c); Dai et al. (2023), our key design is to incorporate references to bounding boxes into language instructions, thereby upgrading them to the format of spatial instructions. For example, as shown in Figure 1, when the question is “what is <region1> doing?”, where the <region1> refers to a specific region-of-interest, the model will substitute the embedding of <region1> with the region feature extracted by the corresponding bounding box. The region feature extractor can be flexibly implemented by RoIAlign He et al. (2017) or Deformable attention Zhu et al. (2020).
@@
To establish fine-grained alignment between vision and language, we involve region-text datasets in our training, where the bounding box and the text description of each region are provided. The datasets are consolidated from publicly available ones including COCO object detection Lin et al. (2014), RefCOCO Yu et al. (2016), RefCOCO+ Yu et al. (2016), RefCOCOg Mao et al. (2016), Flickr30K entities Plummer et al. (2015), Visual Genome (VG) Krishna et al. (2017) and Visual Commonsense Reasoning (VCR) Zellers et al. (2019a). These datasets are transformed into spatial instruction tuning format. Moreover, we incorporate the LLaVA150K dataset Liu et al. (2023a) into our training process by utilizing an off-the-shelf detector to generate bounding boxes. This enhances our model’s ability to engage in multi-round conversations and generate more human-like responses.
@@
Benefiting from spatial instruction tuning, our model brings a new interactive experience, where the user can express the question to the model with language and the reference to the region-of-interest. This leads to new capacities beyond image-level understanding, such as region caption and complex region reasoning. As a generalist, our model GPT4RoI also shows its strong region understanding ability on popular benchmarks, including comprehensive image region understanding benchmark ViP-Bench Cai et al. (2023), the region caption task on Visual Genome Krishna et al. (2017), the region reasoning task on Visual-7W Zhu et al. (2016) and Visual Commonsense Reasoning Zellers et al. (2019a) (VCR). Especially noteworthy is the performance on the most challenging VCR dataset, where GPT4RoI achieves an impressive accuracy of 81.6%, 6 points ahead of the second-place and nearing the human-level performance benchmarked at 85.0%.
@@ 2.1 Large Language Model
The field of natural language processing (NLP) has achieved significant development by the high-capability large language model (LLM). The potential of LLM is first demonstrated by pioneering works such as BERT Devlin et al. (2018) and GPT Radford et al. (2018). Then scaling up progress is started and leads to a series of excellent works, for example, T5 Raffel et al. (2020), GPT-3 Brown et al. (2020), Flan-T5 Chung et al. (2022), PaLM Chowdhery et al. (2022), etc. With the growth of training data and model parameters, this scaling-up progress brings a phenomenal product, ChatGPT OpenAI (2022). By generative pre-trained LLM and instruction tuning Ouyang et al. (2022) on human feedback, ChatGPT shows unprecedented performance on conversations with humans, reasoning and planning tasks Mu et al. (2023); Yang et al. (2023a); Bubeck et al. (2023), etc.
@@ 2.2 Large Vision-Language Model
To utilize high-performance LLM to build up vision-language models, LLM as task coordinator is proposed. Given the user instruction, LLM parses the instruction and calls various external vision models. Some representative works are Visual ChatGPT Wu et al. (2023), ViperGPT Surís et al. (2023), MM-REACT Yang et al. (2023c), InternGPT Liu et al. (2023d), VideoChat Li et al. (2023b), etc. Although these models largely expand the scope of multimodal models, they depend on external vision models and these non-end-to-end architectures are not the optimal choice for multi-modal models. To obtain end-to-end vision-language models, instruction tuning LLM on image-text pairs is proposed to align visual features with LLM and accomplish multimodal tasks in a unified way, for example, Flamingo Alayrac et al. (2022), MiniGPT-4 Zhu et al. (2023), LLaVA Liu et al. (2023a), LLaMa-Adapter Zhang et al. (2023c), InstructBLIP Dai et al. (2023), MM-GPT Gong et al. (2023), VPGTrans Zhang et al. (2023a), etc. These models achieve amazing image-level multimodal abilities, while several benchmarks such as LVLM-eHub Xu et al. (2023) and MMBench Liu et al. (2023c) find that these models still have performance bottlenecks when need to be under specific region reference. Our GPT4RoI follows the research line of visual instruction tuning and moves forward region-level multimodal understanding tasks such as region caption Krishna et al. (2017) and reasoning Zellers et al. (2019a).
@@ 2.3 Region-Level Image Understanding
For region-level understanding, it is a common practice in computer vision to identify potential regions of interest first and then do the understanding. Object detection Ren et al. (2015); Carion et al. (2020); Zhu et al. (2020); Zang et al. (2023) tackles the search for potential regions, which are generally accompanied by a simple classification task to understand the region’s content. To expand the object categories,  Kamath et al. (2021); Liu et al. (2023b); Zhou et al. (2022); Li* et al. (2022) learn from natural language and achieve amazing open-vocabulary object recognition performance. Region captioning Johnson et al. (2015); Yang et al. (2017); Wu et al. (2022) provides more descriptive language descriptions in a generative way. Scene graph generation Li et al. (2017); Tang et al. (2018); Yang et al. (2022) analyzes the relationships between regions by the graph. The VCR Zellers et al. (2019b) dataset presents many region-level reasoning cases and  Yu et al. (2021); Su et al. (2019); Li et al. (2019b); Yao et al. (2022) exhibit decent performance by correctly selecting the answers in the multiple-choice format. However, a general-purpose region understanding model has yet to emerge. In this paper, by harnessing the powerful large language model Touvron et al. (2023a); Chiang et al. (2023), GPT4RoI uses a generative approach to handle all these tasks. Users can complete various region-level understanding tasks by freely asking questions.
@@
We also compare the design philosophy with concurrent works that utilize textual coordinates or extra position tokens as region references, all of which can be found in the Appendix A.4.
@@ 3 GPT4RoI
Figure 2: GPT4RoI is an end-to-end vision-language model for processing spatial instructions that contain references to the region-of-interest, such as <region{i}>. During tokenization and conversion to embeddings, the embedding of <region{i}> in the instruction is replaced with the RoIAlign results from multi-level image features. Subsequently, such an interleaved region feature and language embedding sequence can be sent to a large language model (LLM) for further processing. We also utilize the entire image feature to capture global information and omit it in the figure for brevity. A more detailed framework figure can be found in Figure 4 in the Appendix.
GPT4RoI consists of a vision encoder, a projector for image-level features, a region feature extractor, and a large language model (LLM). Compared to previous works Zhu et al. (2023); Liu et al. (2023a), GPT4RoI stands out for its ability to convert instructions that include spatial positions into an interleaved sequence of region features and text embeddings, as shown in Figure 2.
@@ 3.1 Model Architecture
We adopt the CLIP ViT-L/14 Radford et al. (2021) as the vision encoder. We use the feature map of the penultimate transformer layer as the representation of the entire image and then map it to the language space using a single linear layer. A region feature extractor is used to obtain region features with the region references in the user instruction. Finally, we employ the Vicuna Zheng et al. (2023), an instruction-tuned LLaMA Touvron et al. (2023a), to perform further processing.
We utilize widely adopted modules in the field of object detection to construct our region feature extractor. A multi-level feature pyramid Lin et al. (2017) is constructed based on ViT-L/14. We adopt RoIAlign He et al. (2017) to extract region features on each level with region references in the user instruction. The multi-level region features for each reference are then fused to a single embedding, serving as the region representation(as shown in  2). More details of this region feature extractor can be found in the Appendix. A.1.
@@ 3.3 Spatial Instruction Tuning
Our model is trained using a next-token prediction loss Liu et al. (2023a); Zhu et al. (2023), where the model predicts the next token in a given input text sequence.
The training details are in the Appendix. A.2.
@@
Stage 1: Pre-training In this stage, we first load the weights of LLaVA Liu et al. (2023a) after its initial stage of training, which includes a pre-trained vision encoder, a projector for image-level features, and an LLM. We only keep the region feature extractor trainable and aim to align region features with language embedding by collecting short text and bounding box pairs. These pairs are from both normal detection datasets and referring expression detection datasets, which have short expressions. The objective is to enable the model to recognize categories and simple attributes of the region in an image, which are typically represented by a short text annotation (usually within 5 words). Specifically, we utilize COCO Lin et al. (2014), RefCOCO Yu et al. (2016), and RefCOCO+ Yu et al. (2016) datasets in this stage.
@@
<region3> right most person [Uncaptioned image]
After this training stage, GPT4RoI can recognize categories, simple attributes, and positions of regions in images, as shown in the Figure 5 of Appendix
@@
As shown in Table 2, we tailor specific instructions for different tasks. For single region caption, we construct from Visual Genome (VG) region caption part Krishna et al. (2017) and RefCOCOg Mao et al. (2016). For multiple region caption, Flicker30k Plummer et al. (2015) is converted to a multiple region caption task where the caption should include all visual elements emphasized by bounding boxes. To simulate user instruction, we create 20 questions for each caption task as shown in Table 9 and Table 10 of the Appendix. For the region reasoning task, we modify Visual Commonsense Reasoning (VCR) Zellers et al. (2019a) to meet the input format requirements and make it more similar to human input. The details of this process can be found in Appendix. A.3.
@@
To improve the capability of GPT4RoI for multi-round conversation and generate more human-like responses, we also involve the LLaVA150k Liu et al. (2023a) visual instruction dataset in this stage. We employ an off-the-shelf LVIS detector Fang et al. (2023) to extract up to 100 detection boxes per image. These boxes are then concatenated with the user instructions in the format “<region{i}> may feature a class_name”. LLaVA150k significantly improves the capability of GPT4RoI for multi-round conversation.
@@ 4 Demostrations
In this section, we compare the differences between the visual instruction tuning model LLaVA Liu et al. (2023a) and our spatial instruction tuning model GPT4RoI. We demonstrate our new interactive approach and highlight its advanced capabilities in understanding multimodality.
@@ 5 Experiments
We adopt several representative benchmarks to quantitatively evaluate the region understanding ability of GPT4RoI. First, we give the results on comprehensive LLM image region understanding benchmark ViP-Bench Cai et al. (2023). We also evaluate the region recognition ability through the open-vocabulary region recognition and region caption task. For region reasoning ability, we report the results on Visual-7W Zhu et al. (2016), and Visual Commonsense Reasoning Zellers et al. (2019a).
@@ 5.1 Comprehensive Region Understanding
ViP-Bench Cai et al. (2023) is a comprehensive benchmark for evaluating multimodal models’ region understanding capabilities, including recognition (Rec), OCR, knowledge (Know), math, relationship (Rel), and language generation (Lang). As shown in Table 3, we give the comparison of methods including InstructBLIP Dai et al. (2023) with a pure language reference, Shikra  Chen et al. (2023b) with textual coordinates as reference, Kosmos-2  Peng et al. (2023) with extra position tokens. We can see that GPT4RoI surpasses by a clear margin with actually much less training data.
Type Init. Date Rec OCR Know Math Rel Lang All
Kosmos-2 Jun. 2023 29.5 14.2 18.5 9.7 7.5 21.9 26.9
InstructBLIP May 2023 36.9 16.3 34.2 22.3 26.8 7.5 31.7
Shikra-7B Jul. 2023 40.2 10.0 28.0 3.5 18.9 20.6 33.7
GPT4RoI-7B Jul. 2023 35.6 16.7 29.7 9.7 32.5 13.8 35.1
Table 3: Evaluation results on ViP-Bench.The assessed dimensions are Recognition (Rec), OCR, Knowledge (Know), Math, Relationship (Rel), and Language Generation (Lang). We also provide the initial release date of each work. means concurrent work.

### v5

@@ Abstract
Keywords:
Visual Instruction Tuning LLM Multimodal
@@ 1 Introduction
Figure 1: Comparison of visual instruction tuning on image-text pairs and spatial instruction tuning on region-text pairs. The bounding box and text description of each object are provided in region-text datasets. During training, the bounding box is from annotations, and in inference, it can be provided by the user or any off-the-shelf object detector.
@@
Recent advancements of large language models (LLM) have shown incredible performance in solving natural language processing tasks in a human-like conversational manner, for example, commercial products  [47, 2, 23, 48] and community open-source projects [68, 69, 67, 11, 18, 64]. Their unprecedented capabilities present a promising path toward general-purpose artificial intelligence models. Witnessing the power of LLM, the field of multimodal models [81, 26, 21, 17] is developing a new technology direction to leverage LLM as the universal interface to build general-purpose models, where the feature space of a specific task is tuned to be aligned with the feature space of pre-trained language models.
Table 1: Comparisons of vision-language models. Our GPT4RoI is an end-to-end model that supports region-level understanding and multi-round conversation.
Model Image Region Multi-Region Multi-Round End-to-End
Dialogue Model
[…]
@@
As one of the representative tasks, vision-and-language models align the vision encoder feature to LLM by instruction tuning on image-text pairs, such as MiniGPT-4 [100], LLaVA [37], InstructBLIP [15], etc. Although these works achieve amazing multimodal abilities, their alignments are only on image-text pairs [9, 62, 7, 49, 61], the lack of region-level alignment limits their advancements to more fine-grained understanding tasks such as region caption [29] and reasoning [88]. To enable region-level understanding in vision-language models, some works attempt to leverage external vision models, for example, MM-REACT [81], InternGPT [42] and DetGPT [52], as shown in Table 1. However, their non-end-to-end architecture is a sub-optimal choice for general-purpose multi-modal models.
@@
Considering the limitations of previous works, our objective is to construct an end-to-end vision-language model that supports a fine-grained understanding on region-of-interest. Since there is no operation that can refer to specific regions in current image-level vision-language models [100, 37, 94, 15], our key design is to incorporate references to bounding boxes into language instructions, thereby upgrading them to the format of spatial instructions. For example, as shown in Figure 1, when the question is “what is <region1> doing?”, where the <region1> refers to a specific region-of-interest, the model will substitute the embedding of <region1> with the region feature extracted by the corresponding bounding box. The region feature extractor can be flexibly implemented by RoIAlign [25] or Deformable attention [101].
@@
To establish fine-grained alignment between vision and language, we involve region-text datasets in our training, where the bounding box and the text description of each region are provided. The datasets are consolidated from publicly available ones including COCO object detection [36], RefCOCO [85], RefCOCO+ [85], RefCOCOg [45], Flickr30K entities [54], Visual Genome (VG) [29] and Visual Commonsense Reasoning (VCR) [88]. These datasets are transformed into spatial instruction tuning format. Moreover, we incorporate the LLaVA150K dataset [37] into our training process by utilizing an off-the-shelf detector to generate bounding boxes. This enhances our model’s ability to engage in multi-round conversations and generate more human-like responses.
@@
Benefiting from spatial instruction tuning, our model brings a new interactive experience, where the user can express the question to the model with language and the reference to the region-of-interest. This leads to new capacities beyond image-level understanding, such as region caption and complex region reasoning. As a generalist, our model GPT4RoI also shows its strong region understanding ability on popular benchmarks, including comprehensive image region understanding benchmark ViP-Bench [5], the region caption task on Visual Genome [29], the region reasoning task on Visual Commonsense Reasoning [88] (VCR). Especially noteworthy is the performance on the most challenging VCR dataset, where GPT4RoI achieves an impressive accuracy of 81.6%, 6 points ahead of the second-place and nearing the human-level performance benchmarked at 85.0%.
@@ 2.1 Large Language Model
The field of natural language processing (NLP) has achieved significant development by the high-capability large language model (LLM). The potential of LLM is first demonstrated by pioneering works such as BERT [16] and GPT [57]. Then scaling up progress is started and leads to a series of excellent works, for example, T5 [58], GPT-3 [3], Flan-T5 [13], PaLM [12], etc. With the growth of training data and model parameters, this scaling-up progress brings a phenomenal product, ChatGPT [47]. By generative pre-trained LLM and instruction tuning [50] on human feedback, ChatGPT shows unprecedented performance on conversations with humans, reasoning and planning tasks [46, 78, 4], etc.
@@ 2.2 Large Vision-Language Model
To utilize high-performance LLM to build up vision-language models, LLM as task coordinator is proposed. Given the user instruction, LLM parses the instruction and calls various external vision models. Some representative works are Visual ChatGPT [75], ViperGPT [65], MM-REACT [81], InternGPT [42], VideoChat [31], etc. Although these models largely expand the scope of multimodal models, they depend on external vision models and these non-end-to-end architectures are not the optimal choice for multi-modal models. To obtain end-to-end vision-language models, instruction tuning LLM on image-text pairs is proposed to align visual features with LLM and accomplish multimodal tasks in a unified way, for example, Flamingo [1], MiniGPT-4 [100], LLaVA [37], LLaMa-Adapter [94], InstructBLIP [15], MM-GPT [22], VPGTrans [92], etc. These models achieve amazing image-level multimodal abilities, while several benchmarks such as LVLM-eHub [77] and MMBench [41] find that these models still have performance bottlenecks when need to be under specific region reference. Our GPT4RoI follows the research line of visual instruction tuning and moves forward region-level multimodal understanding tasks such as region caption [29] and reasoning [88].
@@ 2.3 Region-Level Image Understanding
For region-level understanding, it is a common practice in computer vision to identify potential regions of interest first and then do the understanding. Object detection [60, 6, 101, 87] tackles the search for potential regions, which are generally accompanied by a simple classification task to understand the region’s content. To expand the object categories,  [28, 39, 99, 33] learn from natural language and achieve amazing open-vocabulary object recognition performance. Region captioning [27, 80, 76] provides more descriptive language descriptions in a generative way. Scene graph generation [34, 66, 79] analyzes the relationships between regions by the graph. The VCR [89] dataset presents many region-level reasoning cases and  [84, 63, 32, 82] exhibit decent performance by correctly selecting the answers in the multiple-choice format. However, a general-purpose region understanding model has yet to emerge. In this paper, by harnessing the powerful large language model [68, 11], GPT4RoI uses a generative approach to handle all these tasks. Users can complete various region-level understanding tasks by freely asking questions.
@@
Many recent studies (concurrent works or follow-ups of GPT4RoI) integrate region-level data via different region reference formats into the training of MLLMs [51, 8, 96, 53, 55, 83, 86]. In particular, Kosmos-2 [51], Shikra [8], and CogVLM [71] directly quantize bounding boxes into discrete location tokens or numeric representations of positions. Regiongpt [24], Osprey [86], and Groma [44] use different feature extraction operations to extract the features of user reference, which may be partially influenced by our approach. Additionally, there has been research that explores a fusion of the two approaches, as shown in references [83, 59, 93].
@@ 3 GPT4RoI
GPT4RoI consists of a vision encoder, a projector for image-level features, a region feature extractor, and a large language model (LLM). Compared to previous works [100, 37], GPT4RoI stands out for its ability to convert instructions that include spatial positions into an interleaved sequence of region features and text embeddings, as shown in Figure 2.
@@ 3.1 Model Architecture
We adopt the CLIP ViT-L/14 [56] as the vision encoder. We use the feature map of the penultimate transformer layer as the representation of the entire image and then map it to the language space using a single linear layer. A region feature extractor is used to obtain region features with the region references in the user instruction. Finally, we employ the Vicuna [97], an instruction-tuned LLaMA [68], to perform further processing.
Figure 2: GPT4RoI is an end-to-end vision-language model for processing spatial instructions that contain references to the region-of-interest, such as <region{i}>. During tokenization and conversion to embeddings, the embedding of <region{i}> in the instruction is replaced with the RoIAlign results from multi-level image features. Subsequently, such an interleaved region feature and language embedding sequence can be sent to a large language model (LLM) for further processing. We also utilize the entire image feature to capture global information.
We utilize widely adopted modules in the field of object detection to construct the region feature extractor. To ensure a robust feature representation for regions of varying scales, we construct a multi-level image feature pyramid [35] by selecting four layers from the CLIP vision encoder and fusing them with five lightweight scale shuffle modules [95]. These layers are located at the second-to-last, fifth-to-last, eighth-to-last, and eleventh-to-last positions, respectively. Additionally, we incorporate feature coordinates [38, 73] for each level to address the problem of translation invariance in CNNs. This helps make the model sensitive to absolute position information, such as the description “girl on left” in Figure 3. Finally, we use RoIAlign to extract region-level features with an output size of 14$\times$14 [25], which ensures that sufficient detailed information is preserved. Moreover, all four level features are involved in the RoIAlign operation and fused into a single embedding as the representation of the region of interest (RoI) [40].
@@ 3.3 Spatial Instruction Tuning
Our model is trained using a next-token prediction loss [37, 100], where the model predicts the next token in a given input text sequence.
@@
Stage 1: Pre-training In this stage, we first load the weights of LLaVA [37] after its initial stage of training, which includes a pre-trained vision encoder, a projector for image-level features, and an LLM. We only keep the region feature extractor trainable and aim to align region features with language embedding by collecting short text and bounding box pairs. These pairs are from both normal detection datasets and referring expression detection datasets, which have short expressions. The objective is to enable the model to recognize categories and simple attributes of the region in an image, which are typically represented by a short text annotation (usually within 5 words). Specifically, we utilize COCO [36], RefCOCO [85], and RefCOCO+ [85] datasets in this stage.
@@
<region3> right most person
After this training stage, GPT4RoI can recognize categories, simple attributes, and positions of regions in images, as shown in Figure 3.
@@
As shown in Table 3, we tailor specific instructions for different tasks. For single region caption, we construct from Visual Genome (VG) region caption part [29] and RefCOCOg [45]. For multiple region caption, Flicker30k [54] is converted to a multiple region caption task where the caption should include all visual elements emphasized by bounding boxes. To simulate user instruction, we create 20 questions for each caption task. For the region reasoning task, we modify Visual Commonsense Reasoning (VCR) [88] to meet the input format requirements and make it more similar to human input.
@@
To improve the capability of GPT4RoI for multi-round conversation and generate more human-like responses, we also involve the LLaVA150k [37] visual instruction dataset in this stage. We employ an off-the-shelf LVIS detector [19] to extract up to 100 detection boxes per image. These boxes are then concatenated with the user instructions in the format “<region{i}> may feature a class_name”. LLaVA150k significantly improves the capability of GPT4RoI for multi-round conversation.
@@ 4 Demostrations
In this section, we compare the differences between the visual instruction tuning model LLaVA [37] and our spatial instruction tuning model GPT4RoI. We demonstrate our new interactive approach and highlight its advanced capabilities in understanding multimodality.
@@ 5 Experiments
We adopt several representative benchmarks to quantitatively evaluate the region understanding ability of GPT4RoI. First, we give the results on comprehensive LLM image region understanding benchmark ViP-Bench [5]. We also evaluate the region recognition ability through the open-vocabulary region recognition and region caption task. For region reasoning ability, we report the results on Visual Commonsense Reasoning [88].
@@ 5.1 Comprehensive Region Understanding
ViP-Bench [5] is a comprehensive benchmark for evaluating multimodal models’ region understanding capabilities, including recognition (Rec), OCR, knowledge (Know), math, relationship (Rel), and language generation (Lang). As shown in Table 4, we give the comparison of methods including InstructBLIP [15] with a pure language reference, Shikra  [8] with textual coordinates as reference, Kosmos-2  [51] with extra position tokens. We can see that GPT4RoI surpasses by a clear margin with actually much less training data.
Table 4: Evaluation results on ViP-Bench.The assessed dimensions are Recognition (Rec), OCR, Knowledge (Know), Math, Relationship (Rel), and Language Generation (Lang). means concurrent work.
Type Rec OCR Know Math Rel Lang All
Kosmos-2 29.5 14.2 18.5 9.7 7.5 21.9 26.9
InstructBLIP 36.9 16.3 34.2 22.3 26.8 7.5 31.7
Shikra-7B 40.2 10.0 28.0 3.5 18.9 20.6 33.7
GPT4RoI-7B 35.6 16.7 29.7 9.7 32.5 13.8 35.1